OpenStudy (anonymous):

(7x^2+x+1)-(6x^2+x-1)

OpenStudy (anonymous):

just subtract the like terms

OpenStudy (anonymous):

x^2+2

OpenStudy (anonymous):

first, 7x^2-6x^2= x^2

OpenStudy (anonymous):

7x^2-6x^2 = x^2 x-x = 0 1--1 = 1+1= 2

OpenStudy (anonymous):

second, x - (+)x = 0

OpenStudy (anonymous):

third, 1 - (-) 1= 2

OpenStudy (anonymous):

so the answer would be x^2+2